# Bounds on the realizations of zero-nonzero patterns and sign conditions of polynomials restricted to varieties and applications

### v2

@@ Abstract.
We give several applications of our results. We generalize existing results on bounding the speeds of algebraically defined classes of graphs, bounding the ${\varepsilon}$-entropy of real varieties, as well as lower bounds in terms of the number of connected components for testing membership in semi-algebraic sets using algebraic computation trees. Motivated by quantum complexity theory we introduce a notion of relative rank (additive as well as multiplicative) in finite dimensional vector spaces and algebras relative to a fixed algebraic subset in the vector space or algebra – which generalizes the classical definition of ranks of tensors. We prove a very general lower bound on the maximum relative rank of finite subsets relative to algebraic subsets of bounded degree and dimension which is independent of the dimension of the vector space or algebra. We show how our lower bound implies a quantum analog of a classical lower bound result of Shannon for Boolean circuits – that almost all Boolean functions require (classical) circuits of size at least $\Omega(2^{n}/n)$ even in the presence of a quantum oracle given by an algebraic subset (of the group of unitary transformations) of fixed degree and dimension.
@@ 1.2. Basic definitions and notation
where $\mbox{\rm H}_{i}(S,\mathbb{Z}/2\mathbb{Z})$ denotes the $i$-th homology group of $S$ with coefficients in $\mathbb{Z}/2\mathbb{Z}$ (see for example, [7, Chapter 6, Section 3]) and call $b_{i}(S)$ the $i$-th Betti number of $S$ (with coefficients in $\mathbb{Z}/2\mathbb{Z}$). Note that $b_{0}(S)$ is just the number of semi-algebraically connected components of $S$.  [^2] We say “semi-algebraically connected” rather than connected since $\mathrm{R}$ is assumed to be an arbitrary real closed field, possibly non-archimedean. “Semi-algebraically connected” means “semi-algebraically path connected” (see for example [7, Chapter 3, Section 2]), i.e. every two points in the set can be joined by a semi-algebraic path whose image is contained in the set. If $\mathrm{R}=\mathbb{R}$, then being semi-algebraically connected and being connected are equivalent properties of semi-algebraic sets.
@@
[4, Theorem 4] Suppose that
@@
[47, Theorem 2.12] Let $\mathrm{R}$ be a real closed field, $\mathrm{C}=\mathrm{R}[i]$ the algebraic closure of $\mathrm{R}$, and $V\subset\mathrm{C}^{N}$ be an algebraic variety with $\dim V=p$, and $P\in\mathrm{R}[X_{1},\ldots,X_{n}]$. Then the number of semi-algebraically connected components $C$ of $\mathrm{R}^{N}-{\rm Z}(P,\mathrm{R}^{N}))$ such that $C\cap V(\mathrm{R})\neq\emptyset$ is bounded by $C(N)\cdot\deg(V)\cdot\deg(P)^{p}$ where the constant $C(N)$ depends on $N$ (in an unspecified way), and $V(\mathrm{R})$ denotes the $\mathrm{R}$-points of $V$.
@@
[31, Theorem 1.2] The sum of all $\mathbb{Z}/2\mathbb{Z}$-Betti numbers of $V(\mathrm{R})$, where $V$ is a non-singular real variety of degree $D$ and dimension $p$, is bounded by
@@ 1.3. Summary of main contributions
The independence of these bounds from the dimension $N$ of the ambient space is a new feature of these bounds. This feature is potentially useful in any application where one is interested in the number of zero-nonzero patterns (resp. sign conditions) restricted to algebraic subsets of dimension much smaller than that of the ambient dimension. We discuss examples of such applications below in Section 1.3.4.
@@ 1.3.1. Methods
Our techniques for dealing with zero-nonzero patterns and sign conditions are different. To prove the upper bound on the number of zero-nonzero patterns of a family of polynomials restricted to an algebraic subset $V$ of $\mathbb{A}_{k}^{N}$, we closely follow the proof given in [25, 8], with the modification that all algebraic subsets that are considered in the proof are contained in $V$ and we can bound their degrees in terms of $\deg V$ and the degrees of the given polynomials. We prove an upper bound on the sum of the degrees of the Zariski closure of the realizations of all zero-nonzero patterns with non-empty realizations on $V$, which in turn gives an upper bound on the number of realized zero-nonzero patterns. An important role in the argument is played by the generalized Bezout inequality (Theorem I).
@@ 1.3.2. Tightness
1.3.3. Connection with model theory and NIP structures
Bounding the number of realizable zero-nonzero patterns of $n$ subsets of a fixed definable set $V$ all coming from fixed definable family plays an important role in model theory in relation to an important class of theories called NIP theories. These are complete theories whose models have the property that every definable family of subsets has finite Vapnik-Chervonenkis dimension (see for example [35, Chapter 12, Section 4] or [42] for the precise definition of NIP theories and structures). A typical result (valid for instance over o-minimal theories, theories of algebraically closed fields, algebraically closed valued fields, which are all NIP) is that the number of realized zero-nonzero patterns on a fixed definable set $V$ realized by $n$ definable sets drawn from a definable family $\mathcal{H}$ is bounded by
$C_{V,\mathcal{H}}\cdot n^{\dim V},$
where $C_{V,\mathcal{H}}>0$ is a constant depending only on $V$ and $\mathcal{H}$ (see for example [3, 2, 5, 10] for such bounds in different theories). The exponent $\dim V$ is then an upper bound on the Vapnik-Chervonenkis codensity (see any of the papers mentioned above for definition) of the family $\mathcal{H}$. In fact, if $V$ belongs to some definable family $\mathcal{V}$, the constant, $C_{V,\mathcal{H}}$ can be made to depend just on $\mathcal{V},\mathcal{H}$. If we formulate Theorem 1 in the above language, the corresponding families $\mathcal{H}$ and $\mathcal{V}$ are as follows:
$\mathcal{H}=\mathcal{H}_{d}:=\bigcup_{N>0}\mathcal{H}_{d,N},$
where $\mathcal{H}_{d,N}$ consists of hypersurfaces of degree $d$ in $k^{N}$, and
$\mathcal{V}=\mathcal{V}_{D,p}:=\bigcup_{N>0}{V}_{D,p,N}$
where ${V}_{D,p,N}$ is the family of algebraic subsets of $k^{N}$ of degree bounded by $D$ and dimension $p$. Theorem 1 implies that the number of zero-nonzero patterns realized by any $n$ members of $\mathcal{H}$ on any member $V$ of $\mathcal{V}$ is bounded by $C_{\mathcal{V},\mathcal{H}}\cdot n^{\dim V}$, with $C_{\mathcal{V},\mathcal{H}}=C_{D,d,p}(=(O(Dd))^{p})$. This is similar to the prior results in model theory bounding zero-nonzero patterns mentioned above [5, 3, 2, 10], but notice that the families $\mathcal{H}_{d},\mathcal{V}_{D,p}$ are now countable unions of definable families and are not definable themselves. So the above mentioned implication of Theorem 1 cannot be deduced directly from the model theoretic bounds on zero-nonzero patterns (in the theory of algebraically closed fields). A similar remark applies to Theorem 2 as well.
@@ 1.3.4. Applications
We give several applications of our results. Some of these results improve or generalize prior results leveraging our new bounds. These include upper bounds on the “speeds” (see Definitions 3.1 and 3.2) of algebraically defined families graphs (Theorems 8 and 9), the ${\varepsilon}$-entropy of real varieties (Theorem 10), and lower bounds on the depths of algebraic computations trees testing membership in a semi-algebraic set (Theorem 11). In each of these cases the new bounds depend only on the degree and dimension of the relevant algebraic subset and independent of (or in the ${\varepsilon}$-entropy case very weakly dependent on) the ambient dimension. We also introduce a notion of additive (resp. multiplicative) rank of elements of a vector space (resp. algebra) relative to some algebraic subset $\Delta$ (Definitions 3.4 and 3.5). This is motivated from quantum complexity theory, where $\Delta$ can be thought of as an “quantum oracle” (i.e. a set of unitary operators operating on the Hilbert space associated to a $n$-qubit quantum register each having unit cost). We extend the definition of ranks to subsets of the vector space (resp. algebra), and prove lower bounds on the rank of subsets which have some structure (Theorems 12, 13, 14 and 15). As a consequence, finite subsets which are “grid”-like (i.e. has some algebraic structure) must contain elements of relatively large rank. As an application of the rank lower bound theorems we show how to recover Shannon’s lower bound theorem for classical circuits in the case of quantum circuits equipped with an algebraic oracle $\Delta$ (Theorems 16 and 17). In all these applications, the dimension and the degree of $\Delta$ is small compared to the dimension of the ambient vector space or algebra. So it is crucial that the bounds on the number of zero-nonzero patterns or sign conditions that are at the heart of these results be independent of the ambient dimension (for example, this dimension is exponential in the number of qubits in the application to quantum complexity).
@@
    bounding speeds of algebraically defined families of graphs (Section 3.1);
@@
    bounding the ${\varepsilon}$-entropy of real varieties (Section 3.2);
@@ 2.2. Bound on the number of realizable sign conditions
For $x\in V(\mathrm{R})$, we will denote by $\dim_{x}V(\mathrm{R})$ to be the local real dimension (see for example [7, page 191]) of $V(\mathrm{R})$ at $x$ (as a semi-algebraic set). If $V$ is a non-singular variety of dimension $p$, then $\dim_{x}V(\mathrm{R})=p$ for every $x\in V(\mathrm{R})$.
@@
Aside from the other applications that we discuss in the next section Theorem 4 can be used to simplify proofs of certain quantitative results in discrete geometry. For example in [21], the authors prove a bound on the number of geometric permutations induced by $k$-dimensional transversals to a finite family of convex sets in $\mathrm{R}^{d}$. A key role is played by a bound on the number of realizable sign conditions of a certain finite set of polynomials restricted to the Grassmannian $\mathrm{Gr}(k,d)$ (more accurately, the affine Grassmannian but this is unimportant). In order to get a tight bound on this quantity that does not depend on the dimension of the Plücker emdedding the authors utilize a fact from real algebraic geometry – that the Grassmannian varieties, $\mathrm{Gr}(m,N)$ can be embedded semi-algebraically (not scheme theoretically) as an algebraic subset of $\mathrm{R}^{N^{2}}$ [13, Theorem 3.4.4]. This last fact is somewhat remarkable as there is no such affine embedding of the complex Grassmannian (indeed the complex Grassmannian is compact as a topological space, while every complex affine algebraic set of positive dimension is necessarily non-compact). However, using Theorem 4 (noting that the Grassmannians are all non-singular projective varieties) and the known expressions for the degree and the dimension of $\mathrm{Gr}(m,N)$ (see (3.5)) one can obtain a similar bound as in [21, Theorem 2] without resorting to the clever non-standard embedding of the real Grassmannian, and the new bound is a slight improvement over the original result when $k$ is close to $d$.
@@ 3. Applications
3.1. Speeds of algebraic families of graphs
Let $k$ be an algebraically closed field and suppose that $F\subset k^{n}\times k^{n}$ be a fixed constructible relation (i.e. a constructible subset of $k^{n}\times k^{n}$). A labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called an $F$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in(k^{N})^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F.$
We call the family of $F$-graphs for a fixed constructible relation $F$, a constructible family of graphs.
Now suppose that $\mathrm{R}$ is a real closed field and suppose that $F\subset\mathrm{R}^{n}\times\mathrm{R}^{n}$ be a fixed semi-algebraic relation (i.e. a semi-algebraic subset of $\mathrm{R}^{n}\times\mathrm{R}^{n}$). A labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called an $F$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in(\mathrm{R}^{N})^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F.$
We call the family of $F$-graphs for a fixed semi-algebraic relation $F$, a semi-algebraic family of graphs.
Families of $F$-graphs, where $F$ is a semi-algebraic relation, have been studied in graph theory from several perspectives such as Ramsey theory, Zarankiewicz problem etc. (see for example, [1, 17]). We will be concerned with another very well-studied property, called “speed”, of such families.
The function $G_{F}:\mathbb{N}\rightarrow\mathbb{N}$, defined by
$G_{F}(n)=\mbox{ number of distinct $F$-graphs with vertex set $[1,n]$}$
is called the speed of the family of graphs defined by $F$.
The following theorem (slightly simplified) is a key result in the theory of semi-algebraic graphs.
Theorem G.
[38, Theorem 1.2] For every semi-algebraic relation $F\subset\mathrm{R}^{N}\times\mathrm{R}^{N}$,
$G_{F}(n)\leq n^{(1+o_{F}(1))Nn}$
(where the $o_{F}(1)$ term depends on $F$).  [^3] Theorem 1.2 in [38] is slightly more general in that it gives a bound on the number of edge-labellings taking values in some fixed finite set $\Lambda$. The Theorem stated here is the special case when $\Lambda=\{0,1\}$.
Now suppose that $k$ is an algebraically closed field. Let $F\subset k^{N}\times k^{N})_{n>0}$ (resp. $F\subset\mathbb{P}_{k}^{N}\times\mathbb{P}_{k}^{N}$) be a constructible subset, and $Z\subset k^{N}$ (resp. $Z\subset\mathbb{P}_{k}^{N}$ ) an algebraic subset. We will call a labelled (directed) graph $G=(V,E\subset V\times V)$, with $V=[1,n]:=\{1,\ldots,n\}$ is called a $(Z,F)$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in Z^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F,$
and denote the set of $(Z,F)$-graphs with vertex set $[1,n]$ by $\mathcal{G}_{Z,F}(n)$.
Definition 3.1.
Let $\mathbf{F}=(F_{n}\subset k^{N_{n}}\times k^{N_{n}})_{n>0}$ (resp. $\mathbf{F}=(F_{n}\subset\mathbb{P}_{k}^{N_{n}}\times\mathbb{P}_{k}^{N_{n}})_{n>0}$) a sequence of constructible subsets, and $\mathbf{Z}=(Z_{n}\subset k^{N_{n}})_{n>0}$ (resp. $\mathbf{Z}=(Z_{n}\subset\mathbb{P}_{k}^{N_{n}})_{n>0}$) a sequence of algebraic subsets.
We denote
$G_{\mathbf{Z},\mathbf{F}}(n)=\mathrm{card}(\mathcal{G}_{Z_{n},F_{n}}(n)).$
(Note that if $\mathbf{F}=(F_{n}=F\subset k^{N}\times k^{N})_{n>0}$, and $\mathbf{Z}=(Z_{n}=k^{N})_{n>0}$, then $G_{\mathbf{Z},\mathbf{F}}=G_{F}$).
We prove (with the same notation as above):
Theorem 8.
Let $c_{1},c_{2},c_{3},p>0$ be fixed constants, $\mathbf{Z},\mathbf{F}$ be as in Definition 3.1, and moreover suppose that each $F_{n}$ is a $\mathcal{P}_{n}$-constructible subset for some finite set of polynomials $\mathcal{P}_{n}$ in $2N_{n}$-variables, with $\mathrm{card}(\mathcal{P})\leq c_{1}$, $\deg(P)\leq c_{2}n$ for all $P\in\mathcal{P}_{n}$, $\deg(Z_{n})\leq c_{3}n$, and $\dim Z_{n}\leq p$. Then,
$G_{\mathbf{Z},\mathbf{F}}(n)\leq n^{(1+o(1))(p+1)n}.$
Example 3.1.
For example, we can take (as a randomly chosen example satisfying the hypothesis of Theorem 8) $k=\mathbb{C}$, $N_{n}=n$, $Z_{n}\subset\mathbb{P}_{\mathbb{C}}^{n}$ a $K3$-surface in $\mathbb{P}_{\mathrm{C}}^{n}$ with $\deg(Z_{n})=2n-2$ (there is a $19$-dimensional family of these [22, page 593]), and $F_{n}\subset\mathbb{P}_{\mathbb{C}}^{n}\times\mathbb{P}_{\mathbb{\mathrm{C}}}^{n}$ defined by $H_{n}=0$, where $H_{n}=\sum_{0\leq i,j\leq n}X_{i}^{n}Y_{j}^{n}$. In this example, $\dim Z_{n}=2$, $\deg Z_{n}=2n-2$, and $F_{n}$ is $\{H_{n}\}$-constructible with $\deg(H_{n})=2n$. So using the notation of Theorem 8, $c_{1}=1$, $c_{2}=c_{3}=2$, and $p=2$.
Now suppose that $\mathrm{R}$ is a real closed field.
Definition 3.2.
Let $\mathbf{F}=(F_{n}\subset k^{N_{n}}\times k^{N_{n}})_{n>0}$ (resp. $\mathbf{F}=(F_{n}\subset\mathbb{P}_{\mathrm{R}}^{N_{n}}\times\mathbb{P}_{\mathrm{R}}^{N_{n}})_{n>0}$) a sequence of semi-algebraic subsets, and $\mathbf{Z}=(Z_{n}\subset\mathrm{R}^{N_{n}})_{n>0}$ (resp. $\mathbf{Z}=(Z_{n}\subset\mathbb{P}_{\mathrm{R}}^{N_{n}})_{n>0}$) a sequence of algebraic subsets.
As before, we will call a labelled (directed) graph $(V,E\subset V\times V)$, with $V=\{1,\ldots,n\}$ is called a $(\mathbf{Z},\mathbf{F})$-graph if there exists $(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\in Z_{n}^{n}$, such that
$(i,j)\in E\Leftrightarrow(\mathbf{y}_{i},\mathbf{y}_{j})\in F_{n}.$
The function $G_{\mathbf{Z},\mathbf{F}}:\mathbb{N}\rightarrow\mathbb{N}$ is defined as before.
We prove (with the same notation as above):
Theorem 9.
Let $c_{1},c_{2},c_{3},p>0$ be fixed constants, $\mathbf{Z},\mathbf{F}$ be as in Definition 3.2, and moreover suppose that each $F_{n}$ is a $\mathcal{P}_{n}$-semi-algebraic subset for some finite set of polynomials $\mathcal{P}_{n}$ in $2N_{n}$-variables, with $\mathrm{card}(\mathcal{P})\leq c_{1}$, $\deg(P)\leq c_{2}n$ for all $P\in\mathcal{P}_{n}$, and suppose that each $Z_{n}$ is a finite union of real complete intersection varieties, with $\deg(Z_{n})\leq c_{3}n$, and $\dim Z_{n}\leq p$. Then,
$G_{\mathbf{Z},\mathbf{F}}(n)\leq n^{O(pn)}.$
Remark 12.
Note that Theorems 8 and 9 generalizes Theorem G in several different ways. Unlike in Theorem G, the embedding dimension of the underlying constructible/semi-algebraic family of graphs, as well as the degrees of the polynomials and the algebraic subsets from which the points representing the vertices are chosen are all allowed to grow with $n$ (and the embedding dimension $N_{n}$ is allowed to depend arbitrarily on $n$ and be unbounded as a function of $n$). Because of this flexibility, Theorems 8 and 9 cannot be deduced (unlike Theorem G) from prior results on bounding the number of zero-nonzero patterns (Theorem A) or sign conditions (Theorems B and C).
A particular case might be of interest in the theory of geometrically defined graphs.
Example 3.2 (Schubert graphs).
Let $Z=\mathrm{Gr}(m,N)\subset\mathbb{P}_{k}^{\binom{N}{m}-1}$ denote the Plücker embedding of the Grassmannian of $m$-dimensional subspaces of $k^{N}$, and let $F$ be any constructible relation in $\mathbb{P}_{k}^{\binom{N}{m}-1}\times\mathbb{P}_{k}^{\binom{N}{m}-1}$. For example, the restriction of $F$ to $Z\times Z$ could be a “Schubert-type” relation – namely $(p,q)\in F$ if and only $\dim p\cap q\geq i$ for some fixed $i\geq 0$.
Then,
(3.1) $G_{Z,F}(n)=n^{(1+o(1))(m(N-m)+1)n}.$
Remark 13.
Notice that in Example 3.2 if one applies directly Theorem G then one would obtain
(3.2) $G_{Z,F}(n)\leq n^{(1+o(1))(\binom{N}{m}-1)n}$
which is considerably worse in the exponent for $m\sim N/2$. However, representing $m$-dimensional subspaces in $k^{N}$, using $N\times m$ matrices of full rank, it is possible to apply the same techniques as in the proof of Theorem G to get
(3.3) $G_{Z,F}(n)\leq n^{(1+o(1))mNn}$
which is better than inequality (3.2) but still worse than (3.1).

### v3

@@ Abstract.
We give several applications of our results. We generalize existing results on bounding the ${\varepsilon}$-entropy of real varieties, as well as lower bounds in terms of the number of connected components for testing membership in semi-algebraic sets using algebraic computation trees. Motivated by quantum complexity theory we introduce a notion of relative rank (additive as well as multiplicative) in finite dimensional vector spaces and algebras relative to a fixed algebraic subset in the vector space or algebra – which generalizes the classical definition of ranks of tensors. We prove a very general lower bound on the maximum relative rank of finite subsets relative to algebraic subsets of bounded degree and dimension which is independent of the dimension of the vector space or algebra. We show how our lower bound implies a quantum analog of a classical lower bound result of Shannon for Boolean circuits – that almost all Boolean functions require (classical) circuits of size at least $\Omega(2^{n}/n)$ even in the presence of a quantum oracle given by an algebraic subset of fixed degree and dimension.
@@ 1.2. Basic definitions and notation
where $\mbox{\rm H}_{i}(S,\mathbb{Z}/2\mathbb{Z})$ denotes the $i$-th homology group of $S$ with coefficients in $\mathbb{Z}/2\mathbb{Z}$ (see for example, [5, Chapter 6, Section 3]) and call $b_{i}(S)$ the $i$-th Betti number of $S$ (with coefficients in $\mathbb{Z}/2\mathbb{Z}$). Note that $b_{0}(S)$ is just the number of semi-algebraically connected components of $S$.  [^2] We say “semi-algebraically connected” rather than connected since $\mathrm{R}$ is assumed to be an arbitrary real closed field, possibly non-archimedean. “Semi-algebraically connected” means “semi-algebraically path connected” (see for example [5, Chapter 3, Section 2]), i.e. every two points in the set can be joined by a semi-algebraic path whose image is contained in the set. If $\mathrm{R}=\mathbb{R}$, then being semi-algebraically connected and being connected are equivalent properties of semi-algebraic sets.
@@
[3, Theorem 4] Suppose that
@@
[43, Theorem 2.12] Let $\mathrm{R}$ be a real closed field, $\mathrm{C}=\mathrm{R}[i]$ the algebraic closure of $\mathrm{R}$, and $V\subset\mathrm{C}^{N}$ be an algebraic variety with $\dim V=p$, and $P\in\mathrm{R}[X_{1},\ldots,X_{n}]$. Then the number of semi-algebraically connected components $C$ of $\mathrm{R}^{N}-{\rm Z}(P,\mathrm{R}^{N}))$ such that $C\cap V(\mathrm{R})\neq\emptyset$ is bounded by $C(N)\cdot\deg(V)\cdot\deg(P)^{p}$ where the constant $C(N)$ depends on $N$ (in an unspecified way), and $V(\mathrm{R})$ denotes the $\mathrm{R}$-points of $V$.
@@
[28, Theorem 1.2] The sum of all $\mathbb{Z}/2\mathbb{Z}$-Betti numbers of $V(\mathrm{R})$, where $V$ is a non-singular real variety of degree $D$ and dimension $p$, is bounded by
@@ 1.3. Summary of main contributions
The independence of these bounds from the dimension $N$ of the ambient space is a new feature of these bounds (see Remark 9 about the existence of such bounds). This feature is potentially useful in any application where one is interested in the number of zero-nonzero patterns (resp. sign conditions) restricted to algebraic subsets of dimension much smaller than that of the ambient dimension. We discuss examples of such applications below in Section 1.3.3.
@@ 1.3.1. Methods
Our techniques for dealing with zero-nonzero patterns and sign conditions are different. To prove the upper bound on the number of zero-nonzero patterns of a family of polynomials restricted to an algebraic subset $V$ of $\mathbb{A}_{k}^{N}$, we closely follow the proof given in [21, 6], with the modification that all algebraic subsets that are considered in the proof are contained in $V$ and we can bound their degrees in terms of $\deg V$ and the degrees of the given polynomials. We prove an upper bound on the sum of the degrees of the Zariski closure of the realizations of all zero-nonzero patterns with non-empty realizations on $V$, which in turn gives an upper bound on the number of realized zero-nonzero patterns. An important role in the argument is played by the generalized Bezout inequality (Theorem H).
@@ 1.3.3. Applications
We give several applications of our results. Some of these results improve or generalize prior results leveraging our new bounds. These include upper bounds on the ${\varepsilon}$-entropy of real varieties (Theorem 8), and lower bounds on the depths of algebraic computations trees testing membership in a semi-algebraic set (Theorem 9). In each of these cases the new bounds depend only on the degree and dimension of the relevant algebraic subset and independent of (or in the ${\varepsilon}$-entropy case very weakly dependent on) the ambient dimension. We also introduce a notion of additive (resp. multiplicative) rank of elements of a vector space (resp. algebra) relative to some algebraic subset $\Delta$ (Definitions 3.2 and 3.3). This is motivated from quantum complexity theory, where $\Delta$ can be thought of as an “quantum oracle” (i.e. a set of unitary operators operating on the Hilbert space associated to a $n$-qubit quantum register each having unit cost). We extend the definition of ranks to subsets of the vector space (resp. algebra), and prove lower bounds on the rank of subsets which have some structure (Theorems 10, 11, 12 and 13). As a consequence, finite subsets which are “grid”-like (i.e. has some algebraic structure) must contain elements of relatively large rank. As an application of the rank lower bound theorems we show how to recover Shannon’s lower bound theorem for classical circuits in the case of quantum circuits equipped with an algebraic oracle $\Delta$ (Theorems 14 and 15). In all these applications, the dimension of $\Delta$ is assumed to be small compared to the dimension of the ambient vector space or algebra. So it is crucial that the bounds on the number of zero-nonzero patterns or sign conditions that are at the heart of these results be independent of the ambient dimension (for example, this dimension is exponential in the number of qubits in the application to quantum complexity).
@@
    bounding the ${\varepsilon}$-entropy of real varieties (Section 3.1);
@@ 2.2. Bound on the number of realizable sign conditions
For $x\in V(\mathrm{R})$, we will denote by $\dim_{x}V(\mathrm{R})$ to be the local real dimension (see for example [5, page 191]) of $V(\mathrm{R})$ at $x$ (as a semi-algebraic set). If $V$ is a non-singular variety of dimension $p$, then $\dim_{x}V(\mathrm{R})=p$ for every $x\in V(\mathrm{R})$.
@@
Note that if $V\subset\mathbb{P}_{\mathrm{C}}^{N}$ is a variety with $\dim V=p$, such that $V$ is not contained in any hyperplane of $\mathbb{P}_{\mathrm{C}}^{N}$. Then $N\leq\deg(V)+p-1$ (see for example [9, Proposition 4.5.6]) .  [^3] We thank Greg Blekherman and Rainer Sinn for pointing out this connection with the theory of minimal varieties. Now suppose $V$ is a real variety. Then $V$ can be cut out by real polynomials of degree at most $D$. Now using the basic inequality on the Betti numbers of real algebraic sets due to Oleĭnik and Petrovskiĭ [31] (see for example [5, Proposition 7.28]), we would obtain the bound:
(2.2) $b_{0}(V(\mathrm{R}))$$\leq$ $O(D)^{N}$
$\leq$ $O(D)^{D+p-1}.$
Then for $D>2$, the bound in inequality (2.2) is worse than the bound in Theorem 3. Similar remarks apply to Theorems 2, 4 and 7 as well.
Remark 10.
@@
Aside from the other applications that we discuss in the next section Theorem 4 can be used to simplify proofs of certain quantitative results in discrete geometry. For example in [18], the authors prove a bound on the number of geometric permutations induced by $k$-dimensional transversals to a finite family of convex sets in $\mathrm{R}^{d}$. A key role is played by a bound on the number of realizable sign conditions of a certain finite set of polynomials restricted to the Grassmannian $\mathrm{Gr}(k,d)$ (more accurately, the affine Grassmannian but this is unimportant). In order to get a tight bound on this quantity that does not depend on the dimension of the Plücker emdedding the authors utilize a fact from real algebraic geometry – that the Grassmannian varieties, $\mathrm{Gr}(m,N)$ can be embedded semi-algebraically (not scheme theoretically) as an algebraic subset of $\mathrm{R}^{N^{2}}$ [11, Theorem 3.4.4]. This last fact is somewhat remarkable as there is no such affine embedding of the complex Grassmannian (indeed the complex Grassmannian is compact as a topological space, while every complex affine algebraic set of positive dimension is necessarily non-compact). However, using Theorem 4 (noting that the Grassmannians are all non-singular projective varieties) and the known expressions for the degree and the dimension of $\mathrm{Gr}(m,N)$ (see (3.2)) one can obtain a similar bound as in [18, Theorem 2] without resorting to the clever non-standard embedding of the real Grassmannian, and the new bound is a slight improvement over the original result when $k$ is close to $d$.